Circular No. 6022

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) P. W. Chodas and D. K. Yeomans, Jet Propulsion Laboratory, provide

the following “final” predicted jovian impact times: nucleus A = 21, July 16.833 UT; B = 20, 17.121; C = 19, 17.293; D = 18, 17.491; E = 17, 17.629; F = 16, 18.020; G = 15, 18.311; H = 14, 18.810; K = 12, 19.430; L = 11, 19.923; N = 9, 20.431; P = 8 (= P_{2} = 8b), 20.633; Q = 7 (= Q_{1} = 7a), 20.833; R = 6, 21.226; S = 5, 21.632; T = 4, 21.753; U = 3, 21.909; V = 2, 22.178; W = 1, 22.332. The uncertainty is now ∼ ±0.003 day.

M. Kidger, Instituto de Astrofisica de Canarias, reports that multiple infrared detections have been made of the impact on Jupiter of nucleus A = 21, beginning around July 16.845 UT. The observations were made at 2.2 μm with the Nordic Optical Telescope (La Palma), at 2.1 μm with the Carlos Sanchez Telescope and at 2.3 μm with the 3.5-m telescope at Calar Alto. The observations of a “spot” (∼ 1 mag fainter than Jupiter I = Io) at 2.1 and 2.2 μm lasted until about July 16.854, and the 2.3-μm detection continued beyond July 16.875. Kidger stresses that there was no visible detection of the impact in the methane bands near 613 and 892 nm.

Word of the 2.3-μm detection, described as a plume at about the nominal position of the limb, was also received via the SL9 message center at the University of Maryland. It also appears that the plume was observed at 10μm with the 3.6-m telescope at the European Southern Observatory.

Corrigenda. On IAUC 6019, line 7, for 0.1–0.4 mag read 0.1 mag; line 13, for 0.8–0.9 mag read 0.3 mag; line 14, for 0.4 mag read 0.1 mag.

SUPERNOVA 1994V IN MCG–03-38-025 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (I ∼ 17) on a plate taken on July 14 by Q. A. Parker with the U.K. Schmidt Telescope. The supernova is at α = 14^{h}51^{m}52^{s}.34, δ = −17^{o}12′33′′.5 (equinox 1950.0; uncertainty 0′′.3 in α, 0′′.9 in δ), offset 2′′.6 west and 21′′.9 south of the center of MCG–03-38-025 = IRAS 14518– 1712. The supernova was ∼ 0.5 mag brighter on I plates taken with the same telescope by M. Hartley on June 13 and 15, but it does not appear to a limit of R = 22 on a plate by Hartley on March 10 and several earlier J, R and I plates. A star of I ∼ 16 (position end figures 51^{s}.98, 31′′.8), is situated only 5′′ from the supernova.

1994 July 16 Brian G. Marsden